40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Mon Aug 10 02:11:47 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 IMG_20150803_223403975.jpgIMG_20150803_221210816_HDR.jpgIMG_20150803_223420267.jpg
    Reply  Mon Aug 10 11:39:40 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 
       Reply  Mon Aug 10 11:57:17 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 screen_shot.png
          Reply  Mon Aug 10 12:09:49 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 IMG_20150809_215628585_HDR.jpg
             Reply  Fri Aug 28 01:08:14 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ Calibration RFAM_detector_calib.pdfRFAM_detector_calib_spectra.pdf
                Reply  Fri Aug 28 02:14:53 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ 37MHz OCXO AM measurement OCXO_AM_noise.pdf
                   Reply  Tue Sep 8 10:55:31 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ 37MHz OCXO AM measurement 
Message ID: 234     Entry time: Mon Aug 10 12:09:49 2015     In reply to: 233     Reply to this: 237
Author: Koji 
Type: Electronics 
Category: AM Stabilized EOM Driver 
Subject: RF AM Measurement Unit E1500151 

Still suffering from a power supply issue!

I have been tracking the issues I'm having with the RF AM detector board.

I found that the -5V test point did not show -5V but ~+5V! It seemed that this pin was not connected to -5V but was passive.

I removed the RF AM detector board and exposed the power board again. Pin 11 of P3 interboard connector indeed was not connected to TP12 (-5V). What the hell?

As seen in the attached photo, the PCB pattern for the Pin 11 is missing at the label "!?" and not driven. I soldered a piece of wire there and now Pin11 is at -5V.


This fix actually changed several things. Now the bias setting by the rotary switches works.

Setting BIAS1
 [dBm]   [V]

   0    0.585
   2    0.720
   4    0.897
   6    1.12
   8    1.42
  10    1.79
  12    2.25
  14    2.84
  16    3.60
  18    4.56
  20    5.75
  22    7.37

This allows me to elliminate the saturation of MON1 of the first RF AM detector. I can go ahead to the next step for the first channel.

Now the bias feedback of the second detector is also behaving better. Now TP2 is railing.

Still MON2 is saturated. So, the behavior of the peak detectors needs to be reviewed.

Attachment 1: IMG_20150809_215628585_HDR.jpg  310 kB  | Show | Show all
ELOG V3.1.3-